# Large Language Models for Computer-Aided Design: A Survey

### v1

@@ 4.1.2. The PaLM Family (Google)
PaLM (Chowdhery et al., 2023): It is a 540B-parameter model trained on 780B tokens. It introduces several architectural enhancements like SwiGLU activation, multi-query attention, and shared input-output embeddings. PaLM shows state-of-the-art few-shot performance on a wide range of tasks.
PaLM-2 (Anil et al., 2023): PaLM-2 is a smaller (340B-parameter) but more efficient model, pre-trained on 3.6 trillion (T) tokens. It delivers superior reasoning and multilingual capabilities while reducing training and inference costs.
Med-PaLM / Med-PaLM 2 (Singhal et al., 2023, 2025): Med-PaLM is a domain-specific model fine-tuned for medical tasks. Med-PaLM 2 outperforms its predecessor using domain adaptation and ensemble techniques, achieving strong performance across clinical benchmarks and professional exams.
PaLM-E (Driess et al., 2023): It is a multimodal version with 562B parameters (540B from PaLM and 22B from a vision Transformer (Dehghani et al., 2023)). It supports visual question answering, zero-shot multimodal CoT reasoning, few-shot prompting, OCR-free math reasoning, and multi-image reasoning.
U-PaLM (Tay et al., 2023): U-PaLM improves model quality with minimal additional compute, achieving similar or better performance than PaLM at about half the computational budget.
Flan-PaLM / Flan-U-PaLM (Chung et al., 2024): These versions focus on instruction tuning and CoT prompting. Flan-PaLM, with 540B parameters fine-tuned on 1.8K tasks, significantly outperforms the base PaLM model. Flan-U-PaLM further enhances the performance with advanced adaptation techniques like UL2R (Tay et al., 2023).
@@ 4.1.3. The Gemini Series (Google)
Gemini (Team et al., 2023): Gemini demonstrates strong capabilities across text, image, video, and audio understanding, with support for a 32K context length.
Gemini 1.5 (Team et al., 2024a): Gemini 1.5 brings notable advancements, including a new model architecture, the use of mixture-of-experts (MoE), and training on a much larger dataset containing millions of tokens.
Gemini 2.0 (Pichai et al., 2024): Gemini 2.0 is a major update to its predecessor, improving both speed and performance compared to Gemini 1.5.
Gemini 2.5 (Kavukcuoglu, 2025): Gemini 2.5 further enhances reasoning and coding capabilities, incorporating techniques like CoT prompting for more structured reasoning.
However, Gemma (Team et al., 2024b) and Gemma 2 (Team et al., 2024c) are lightweight, open-source versions of Gemini, providing powerful multimodal capabilities for free.
We summarize state-of-the-art closed-source LLMs in Table 1.
Since ChatGPT, released in November 2022, was the first LLM to demonstrate exceptional performance across a variety of tasks, we focus on models released from 2023 onward. Before ChatGPT, LLMs were generally less powerful, but following its release, models have grown powerful enough to compete with ChatGPT, driven by exponential growth in both model size and data. It is worth noticing that we only list representative LLMs in the table. For example, we list Gemini but omit Gemini 1.5 due to limited technical innovations. However, since PaLM and PaLM-2 represent distinct models, both are included.
Table 1. Summary of state-of-the-art closed source LLMs.
@@
GPT-4V (OpenAI, 2023) - - 2023 OpenAI Vision-enabled GPT-4 variant
@@
OpenAI o3-mini (OpenAI, 2025b) 3B - 2025 OpenAI Lightweight GPT-4 variant for mobile use
GPT-4.5 (OpenAI, 2025a) - - 2025 OpenAI Improved emotional intelligence
@@
PPLXa 70B - 2023 Perplexity AI Retrieval-augmented, mobile-ready LLM
Inflection-2.5b 2.5B - 2024 Inflection AI Personal assistant, dialogue-tuned
Claude 3 (Anthropic, 2024) - - 2024 Anthropic Constitutional AI, safety-tuned
Grok 3c - - 2025 xAI Twitter-integrated, humor-oriented chat
@@ 4.2.1. The LLaMA Family (Meta)
LLaMA-2 (Touvron et al., 2023b): LLaMA-2 includes both foundational models and chat-optimized versions, such as LLaMA-2 Chat. While the architecture remains largely the same as LLaMA, LLaMA-2 was trained on 40% more data. LLaMA-2 Chat further improves safety by fine-tuning on safe response samples and incorporating an additional RLHF step.
LLaMA-3 / 3.1 (Grattafiori et al., 2024): LLaMA-3 / 3.1 models are trained on a dataset seven times larger than that of LLaMA-2. LLaMA-3 comes in two sizes: 8B and 70B. With the largest model featuring 405B parameters, LLaMA-3.1 benefits from 15T tokens of training data and a significantly larger context window (128K compared to LLaMA-2’s 8K). These improvements lead to its competitive performance, achieving results comparable to GPT-4 and GPT-4o.
@@
LLaMA-4 (Meta AI, 2025): LLaMA-4 introduces significant advancements by supporting text, images, audio, and video, enabling seamless multimodal reasoning and generation. This is a major shift from its predecessors, LLaMA-2 and LLaMA-3, which were primarily text-based. Leveraging a Mixture of Experts (MoE) architecture, LLaMA-4 efficiently scales to trillions of parameters, activating only a subset of experts during inference for optimal performance. Additionally, it supports an impressive context length of up to 10 million tokens, making it ideal for long-form tasks like multi-document summarization and complex dialogues.
@@
In addition to the core LLaMA models, a variety of instruction-following models have been developed based on LLaMA or LLaMA-2, including: Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), Guanaco (Dettmers et al., 2023), Koala (Geng et al., 2023), Mistral 7B (Jiang et al., 2023), Code LLaMA (Rozière et al., 2023), Gorilla (Patil et al., 2024), Giraffe (Pal et al., 2023), Vigogne (Huang, 2023), Tulu 65B (Wang et al., 2023a), Long LLaMA (Tworkowski et al., 2023), Stable Beluga 2 (Mahan et al., line), Qwen (Bai et al., 2023), among others. Vicuna is particularly popular for its multimodal language modeling, giving rise to models like LLaVA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2024), InstructBLIP (Dai et al., 2023), and PandaGPT (Su et al., 2023).
4.2.2. The DeepSeek Family (DeepSeek)
DeepSeek-Coder (Guo et al., 2024): DeepSeek-Coder is the first model in the DeepSeek family, followed by DeepSeek-LLM (Bi et al., 2024), DeepSeek-MoE (Dai et al., 2024), and DeepSeek-Math (Shao et al., 2024). DeepSeek-LLM investigates the scaling laws for LLMs to identify the optimal model size and training data scale. DeepSeek-MoE is an innovative MoE architecture specially designed towards ultimate expert specialization. DeepSeek-Math is a domain-specific language model that significantly outperforms the mathematical capabilities of open-source models and approaches the performance level of GPT-4 on academic benchmarks.
DeepSeek-VL (Lu et al., 2024): DeepSeek-VL is an open-source Vision-Language Model designed for real-world applications that combine vision and language understanding.
DeepSeek-V2 (Liu et al., 2024a): DeepSeek-V2 uses multi-head latent attention to reduce inference costs by compressing the key-value cache into a latent vector. This method achieves 5.76 times faster inference throughput compared to the previous DeepSeek models.
DeepSeek-V3 (Liu et al., 2024b): DeepSeek-V3 is a stronger MoE model, featuring 671B parameters, with 37B activated per token. It is pre-trained on 14.8T high-quality, diverse tokens and undergoes SFT and RL stages to fully optimize its capabilities.
DeepSeek-R1 (Guo et al., 2025): DeepSeek-R1-Zero is pre-trained with large-scale RL without the initial SFT stage. DeepSeek-R1 follows a multi-stage training process, including cold-start data before RL, and incorporating two RL stages to improve reasoning patterns and align with human preferences, along with two SFT stages to enhance reasoning and non-reasoning abilities.
@@ 4.2.2. The DeepSeek Family (DeepSeek)
LLaMA-4 (Meta AI, 2025) 109B 40T 2025 Meta 10 million context length, multimodal $+$ MoE
@@
DeepSeek-VL (Lu et al., 2024) 7B 2T 2024 DeepSeek Vision-language aligned
@@
DeepSeek-V3 (Liu et al., 2024b) 671B 14.8T 2024 DeepSeek Flagship MoE LLM
@@
Stable Beluga 2 (Mahan et al., line) 70B - 2023 Stability AI Instruction-tuned LLaMA-2
@@
Vicuna (Chiang et al., 2023) 13B - 2023 LMSYS Chatbot fine-tuned on ShareGPT (anon8231489123, 2023)
@@
StartCoder (Li et al., 2023a) 15.5 1T 2023 - Code generation LLM
@@
CodeT5$+$ (Wang et al., 2023c) 16B 51.5B 2023 - Code understanding $+$ generation
@@
Jamba (Lieber et al., 2024) 52B - 2024 AI21 Labs Hybrid Transformer $+$ MoE
@@
Command R$+$d 104B - 2024 Cohere RAG-enhanced instruction model for finance
@@ 5.1. Data Generation
LLMs have emerged as powerful tools for content generation, making data generation one of their primary applications in the field of CAD. This capability is particularly useful for creating synthetic datasets, fine-tuning models, and addressing the lack of annotated CAD data.
@@
Yuan et al. (Yuan et al., 2024a) utilized GPT-4o (Hurst et al., 2024) to generate datasets that map natural language descriptions and rendered view images to CAD operation sequences. This approach enabled the construction of textual descriptions corresponding to CAD commands. Similarly, Xu et al. (Xu et al., 2025) employed the open-source model InternVL2-26B (Chen et al., 2024b), by randomly selecting four view images of CAD models to generate high-quality textual captions. These captions were used to build a multimodal CAD dataset. Khan et al. (Khan et al., 2024b) adopted a two-step generation process using open-source LLMs and Vision-Language Models (VLMs). First, they rendered multi-view images of individual parts and the final models, which were input to LLaVA-NeXT model (Liu et al., 2024d) with predefined prompts to generate simplified object-level descriptions. Then, they used Mixtral-50B (Jiang et al., 2024) to produce multi-level natural language instructions. Raw CAD sequences from DeepCAD (Wu et al., 2021) were preprocessed to replace meaningless keys with more descriptive terms, which, along with the simplified shape descriptions, served as input to Mixtral-50B. The result was a four-level hierarchy of textual instructions: abstract, simplified, generalized geometric, and detailed geometric descriptions, generated using $k$-shot prompting techniques (Brown et al., 2020).
In another recent work, Wang et al. (Wang et al., 2025b) used a VLM to generate initial draft captions for rendered CAD images, which were then refined through human annotations to create a dataset, pairing these captions with CAD parametric sequences. In (Wang et al., 2025a), Wang et al. employed GPT-4o (Hurst et al., 2024) to classify and filter CAD models, and then used InstructGPT (Wang et al., 2023b) to generate natural language descriptions. The resulting dataset contained CAD modeling sequences aligned with textual descriptions and rendered images from fixed viewpoints.
@@
Other examples of using LLM for CAD data generation include (Du et al., 2024a), which developed BlendNet, a custom training dataset designed to capture diverse communication styles across three axes: 16 object categories, 8 instruction tones, and 5 length categories. Starting with 135 manually crafted seed instructions, the authors used self-instruct data distillation to expand to 50,000 samples. GPT-4o (Hurst et al., 2024) was used to generate scripts from these instructions, which were executed to produce corresponding images. GPT-4o then served as a validator to assess the alignment between images and instructions, resulting in a final dataset of 2,000 human-verified and 6,000 model-validated samples. Similarly, Li et al. (Li et al., 2024a) also addressed the challenge of unavailable textual descriptions in existing CAD datasets. They applied the multimodal model CoCa (Yu et al., 2022) to generate captions for parametric CAD models in the DeepCAD dataset (Wu et al., 2021), creating a new, richly annotated dataset linking text to CAD models.
@@
It is worth noting that current research in data generation primarily leverages the text generation capabilities of LLMs. To date, in the CAD field, there is no work that directly generates other forms of LLM-based data—such as images, point clouds, or other format of data—to construct new datasets in the CAD domain.

### v2

@@ 4.1.2. The PaLM Family (Google)
PaLM (Chowdhery et al., 2023): It is a 540B-parameter model trained on 780B tokens. It introduces several architectural enhancements like SwiGLU activation, multi-query attention, and shared input-output embeddings. PaLM shows state-of-the-art few-shot performance on a wide range of tasks. PaLM-2 (Anil et al., 2023): PaLM-2 is a smaller (340B-parameter) but more efficient model, pre-trained on 3.6 trillion (T) tokens. It delivers superior reasoning and multilingual capabilities while reducing training and inference costs. Med-PaLM / Med-PaLM 2 (Singhal et al., 2023, 2025): Med-PaLM is a domain-specific model fine-tuned for medical tasks. Med-PaLM 2 outperforms its predecessor using domain adaptation and ensemble techniques, achieving strong performance across clinical benchmarks and professional exams. PaLM-E (Driess et al., 2023): It is a multimodal version with 562B parameters (540B from PaLM and 22B from a vision Transformer (Dehghani et al., 2023)). It supports visual question answering, zero-shot multimodal CoT reasoning, few-shot prompting, OCR-free math reasoning, and multi-image reasoning. U-PaLM (Tay et al., 2023): U-PaLM improves model quality while requiring minimal additional computational resources, achieving similar or better performance than PaLM at about half the computational budget. Flan-PaLM / Flan-U-PaLM (Chung et al., 2024): These versions focus on instruction tuning and CoT prompting. Flan-PaLM, with 540B parameters fine-tuned on 1.8K tasks, significantly outperforms the base PaLM model. Flan-U-PaLM further enhances the performance with advanced adaptation techniques like UL2R (Tay et al., 2023).
@@ 4.1.3. The Gemini Series (Google)
Gemini (Team et al., 2023): Gemini demonstrates strong capabilities across text, image, video, and audio understanding, with support for a 32K context length. Gemini 1.5 (Team et al., 2024a): Gemini 1.5 brings notable advancements, including a new model architecture, the use of mixture-of-experts (MoE), and training on a much larger dataset containing millions of tokens. Gemini 2.0 (Pichai et al., 2024): Gemini 2.0 is a major update to its predecessor, improving both speed and performance compared to Gemini 1.5. Gemini 2.5 (Kavukcuoglu, 2025): Gemini 2.5 further enhances reasoning and coding capabilities, incorporating techniques like CoT prompting for more structured reasoning. Gemini 3 (Google DeepMind, 2025): Gemini 3 is Google DeepMind’s AI model offering further advanced reasoning, creative problem-solving, and reliable multimodal understanding. However, Gemma (Team et al., 2024b) and Gemma 2 (Team et al., 2024c) are lightweight, open-source versions of Gemini, providing powerful multimodal capabilities for free.
We summarize state-of-the-art closed-source LLMs in Table 1. Since ChatGPT, released in November 2022, was the first LLM to demonstrate exceptional performance across a variety of tasks, we focus on models released from 2023 onward. Before ChatGPT, LLMs were generally less powerful, but following its release, models have grown powerful enough to compete with ChatGPT, driven by exponential growth in both model size and data. It is worth noticing that we only list representative LLMs in the table. For example, we list Gemini but omit Gemini 1.5 due to limited technical innovations. However, since PaLM and PaLM-2 represent distinct models, both are included.
Table 1. Summary of state-of-the-art closed-source LLMs.
@@
GPT-4V (OpenAI, 2023b) - - 2023 OpenAI Vision-enabled GPT-4 variant
@@
OpenAI o3-mini (OpenAI, 2025d) 3B - 2025 OpenAI Lightweight GPT-4 variant for mobile use
GPT-4.5 (OpenAI, 2025c) - - 2025 OpenAI Improved emotional intelligence
@@
PPLX 70B - 2023 Perplexity AI Retrieval-augmented, mobile-ready LLM
Inflection-2.5 2.5B - 2024 Inflection AI Personal assistant, dialogue-tuned
Claude 3/3.5/4/4.5 (Anthropic, 2024b) - - 2024 Anthropic Constitutional AI, safety-tuned
Grok 3&4 - - 2025 xAI Twitter-integrated, humor-oriented chat
@@ 4.2.1. The LLaMA Family (Meta)
LLaMA-2 (Touvron et al., 2023b): LLaMA-2 includes both foundational models and chat-optimized versions, such as LLaMA-2 Chat. While the architecture remains largely the same as LLaMA, LLaMA-2 was trained on 40% more data. LLaMA-2 Chat further improves safety by fine-tuning on safe response samples and incorporating an additional RLHF step. LLaMA-3 / 3.1 (Grattafiori et al., 2024): LLaMA-3 / 3.1 models are trained on a dataset seven times larger than that of LLaMA-2. LLaMA-3 comes in two sizes: 8B and 70B. With the largest model featuring 405B parameters, LLaMA-3.1 benefits from 15T tokens of training data and a significantly larger context window (128K compared to LLaMA-2’s 8K). These improvements lead to its competitive performance, achieving results comparable to GPT-4 and GPT-4o. LLaMA-4 (Meta AI, 2025): LLaMA-4 introduces significant advancements by supporting text, images, audio, and video, enabling seamless multimodal reasoning and generation. This is a major shift from its predecessors, LLaMA-2 and LLaMA-3, which were primarily text-based. Leveraging a Mixture of Experts (MoE) architecture, LLaMA-4 efficiently scales to trillions of parameters, activating only a subset of experts during inference for optimal performance. Additionally, it supports an impressive context length of up to 10 million tokens, making it ideal for long-form tasks like multi-document summarization and complex dialogues.
@@
In addition to the core LLaMA models, a variety of instruction-following models have been developed based on LLaMA or LLaMA-2, including: Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), Guanaco (Dettmers et al., 2023), Koala (Geng et al., 2023), Mistral 7B (Jiang et al., 2023), Code LLaMA (Rozière et al., 2023), Gorilla (Patil et al., 2024), Giraffe (Pal et al., 2023), Vigogne (Huang, 2023), Tulu 65B (Wang et al., 2023a), Long LLaMA (Tworkowski et al., 2023), Stable Beluga 2 (Mahan et al., 2023), Qwen (Bai et al., 2023), among others. Vicuna is particularly popular for its multimodal language modeling, giving rise to models like LLaVA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2024), InstructBLIP (Dai et al., 2023), and PandaGPT (Su et al., 2023).
@@
LLaMA-4 (Meta AI, 2025) 109B 40T 2025 Meta 10 million context length, multimodal ++ MoE
@@
DeepSeek-VL (Lu et al., 2024a) 7B 2T 2024 DeepSeek Vision-language aligned
@@
DeepSeek-V3/3.1 (Liu et al., 2024b) 671B 14.8T 2024 DeepSeek Flagship MoE LLM
@@
Stable Beluga 2 (Mahan et al., 2023) 70B - 2023 Stability AI Instruction-tuned LLaMA-2
@@
Vicuna (Chiang et al., 2023) 13B - 2023 LMSYS Chatbot fine-tuned on ShareGPT (Anonymous, 2023)
@@
StartCoder (Li et al., 2023b) 15.5 1T 2023 - Code generation LLM
@@
CodeT5++ (Wang et al., 2023c) 16B 51.5B 2023 - Code understanding ++ generation
@@
Jamba (Lenz et al., 2025) 52B - 2024 AI21 Labs Hybrid Transformer ++ MoE
@@
Command R++d 104B - 2024 Cohere RAG-enhanced instruction model for finance
4.2.2. The DeepSeek Family (DeepSeek)
DeepSeek-Coder (Guo et al., 2024): DeepSeek-Coder is the first model in the DeepSeek family, followed by DeepSeek-LLM (Bi et al., 2024), DeepSeek-MoE (Dai et al., 2024), and DeepSeek-Math (Shao et al., 2024). DeepSeek-LLM investigates the scaling laws for LLMs to identify the optimal model size and training data scale. DeepSeek-MoE is an innovative MoE architecture specially designed towards ultimate expert specialization. DeepSeek-Math is a domain-specific language model that significantly outperforms the mathematical capabilities of open-source models and approaches the performance level of GPT-4 on academic benchmarks. DeepSeek-VL (Lu et al., 2024a): DeepSeek-VL is an open-source Vision-Language Model designed for real-world applications that combine vision and language understanding. DeepSeek-V2 (Liu et al., 2024a): DeepSeek-V2 uses multi-head latent attention to reduce inference costs by compressing the key-value cache into a latent vector. This method achieves 5.76 times faster inference throughput compared to the previous DeepSeek models. DeepSeek-V3 / 3.1 (Liu et al., 2024b): DeepSeek-V3 is a stronger MoE model, featuring 671B parameters, with 37B activated per token. It is pre-trained on 14.8T high-quality, diverse tokens and undergoes SFT and RL stages to fully optimize its capabilities. Compared to DeepSeek-V3, DeepSeek-V3.1 doubles the context window, introduces a hybrid reasoning mode, enhances math, logic, and code capabilities, and improves multimodal performance while maintaining efficient, open deployment. DeepSeek-R1 (Guo et al., 2025): DeepSeek-R1-Zero is pre-trained with large-scale RL without the initial SFT stage. DeepSeek-R1 follows a multi-stage training process, including cold-start data before RL, and incorporating two RL stages to improve reasoning patterns and align with human preferences, along with two SFT stages to enhance reasoning and non-reasoning abilities.
@@ 5.1. Data Generation
LLMs have emerged as powerful tools for content generation, making data generation one of their primary applications in the field of CAD. This capability is particularly useful for creating synthetic datasets, fine-tuning models, and addressing the lack of annotated CAD data. Yuan et al. (Yuan et al., 2024b) utilized GPT-4o (Hurst et al., 2024) to generate datasets that map natural language descriptions and rendered view images to CAD operation sequences. This approach enabled the construction of textual descriptions corresponding to CAD commands. Similarly, Xu et al. (Xu et al., 2025) employed the open-source model InternVL2-26B (Chen et al., 2024b), by randomly selecting four view images of CAD models to generate high-quality textual captions. These captions were used to build a multimodal CAD dataset. Khan et al. (Khan et al., 2024a) adopted a two-step generation process using open-source LLMs and Vision-Language Models (VLMs). First, they rendered multi-view images of individual parts and the final models, which were input to LLaVA-NeXT model (Liu et al., 2024d) with predefined prompts to generate simplified object-level descriptions. Then, they used Mixtral-50B (Jiang et al., 2024) to produce multi-level natural language instructions. Raw CAD sequences from DeepCAD (Wu et al., 2021) were preprocessed to replace meaningless keys with more descriptive terms, which, along with the simplified shape descriptions, served as input to Mixtral-50B. The result was a four-level hierarchy of textual instructions: abstract, simplified, generalized geometric, and detailed geometric descriptions, generated using $k$-shot prompting techniques (Brown et al., 2020).
@@
In another recent work, Wang et al. (Wang et al., 2025a) used a VLM: llava-onevision-qwen2-7b (Li et al., 2024a), to generate initial draft captions for rendered CAD images, which were then refined through human annotations to create a dataset, pairing these captions with CAD parametric sequences. In (Wang et al., 2025b), Wang et al. employed GPT-4o (Hurst et al., 2024) to classify and filter CAD models, and then used InstructGPT (Wang et al., 2023b) to generate natural language descriptions. The resulting dataset contained CAD modeling sequences aligned with textual descriptions and rendered images from fixed viewpoints. Xie et al. (Xie and Ju, 2025) built a dataset by converting minimal JSON from Text2CAD (Khan et al., 2024a) into CadQuery code using Gemini 2.0 Flash (Pichai et al., 2024), with guidance and in-context examples. Code failures triggered self-correction (Madaan et al., 2023). Outputs were rendered via Blender (Community, 2024) and compared to ground truth. Samples were accepted only when both model and human agreed on shape similarity, yielding 170000 annotations. Yuan et al. (Yuan et al., 2025) generated synthetic data by leveraging VLMs. They created editing instructions by summarizing differences between the original and edited CAD models from both visual and sequence modalities, using GPT-4o (Hurst et al., 2024) for the visual modality and LLaMA-3-70B (Grattafiori et al., 2024) for the sequence modality. Other examples of using LLM for CAD data generation include (Du et al., 2024), which developed BlendNet, a custom training dataset designed to capture diverse communication styles across three axes: 16 object categories, 8 instruction tones, and 5 length categories. Starting with 135 manually crafted seed instructions, the authors used self-instruct data distillation to expand to 50000 samples. GPT-4o (Hurst et al., 2024) was used to generate scripts from these instructions, which were executed to produce corresponding images. GPT-4o then served as a validator to assess the alignment between images and instructions, resulting in a final dataset of 2000 human-verified and 6000 model-validated samples. Similarly, Li et al. (Li et al., 2024d) also addressed the challenge of unavailable textual descriptions in existing CAD datasets. They applied the multimodal model CoCa (Yu et al., 2022) to generate captions for parametric CAD models in the DeepCAD dataset (Wu et al., 2021), creating a new, richly annotated dataset linking text to CAD models.
Likewise, in (Lv and Bao, 2025), GPT-4o was employed to select CAD model names, while Claude-3.5-Sonnet was used to further refine, filter, and correct them. GPT-4o was also applied for image recognition of rendered views. DeepSeek-VL2 Small (Wu et al., 2024d) generated comprehensive model descriptions, and DeepSeek-V3 (Liu et al., 2024b) produced succinct explanations of the modeling logic for each component. In the same vein, Rosnitschek et al. (Rosnitschek et al., 2025) prompted Grok 3 to generate diverse multi-turn dialogues for constructing a bespoke dataset. Similarly, Govindarajan et al. (Govindarajan et al., 2026) leveraged the multimodal capabilities of GPT-4.1 to generate concise, natural-sounding, yet geometrically precise descriptions, and constructed a new dataset. Guan et al. (Guan et al., 2025) also prompted the code-generation LLM DeepSeek-V3 (Liu et al., 2024b) to generate multiple candidate CadQuery scripts, selected the candidate with the lowest Chamfer Distance as the ground truth, and constructed a corresponding dataset. Moreover, in (Li et al., 2025d), the authors constructed a new CAD dataset of 40k samples following the Sketch, Sketch-based Feature, and Refinements modeling paradigm, with each sample paired with a textual description generated by GPT-4o.
@@
It is worth noting that current research in data generation primarily leverages the text generation capabilities of LLMs. To date, in the CAD field, there is no work that directly generates other forms of LLM-based data, such as images, point clouds, or other format of data, to construct new datasets in the CAD domain.